### v1

@@ Abstract
It has been proposed in medical decision analysis to express the “first do no harm” principle as an asymmetric utility function in which the loss from killing a patient would count more than the gain from saving a life. Such a utility depends on unrealized potential outcomes, and we show how this yields a paradoxical decision recommendation in a simple hypothetical example involving games of Russian roulette. The problem is resolved if we allow the potential outcomes to be random variables. This leads us to conclude that, if you are interested in this sort of asymmetric utility function, you need to move to the stochastic potential outcome framework. We discuss the implications of the choice of parameterization in this setting.
1 Expressing the “first do no harm” principle as an asymmetric utility function on potential outcomes
@@
Causal inference has traditionally focused on identification of average effects. In recent decades, there has been increasing interest in considering different forms of the average causal effect (Imbens and Angrist, 1994), flexible models for individual effects (Hill, 2011, Wager and Athey, 2018), and understanding how effects vary (Ding et al., 2019).
@@
Individual causal effects can never be precisely estimated (without multiple measurements on the same units and strong assumptions about crossover effects), but they can be defined theoretically as latent variables using the potential-outcomes framework (Neyman, 1923, Rubin, 1974). In the simplest case of binary treatment $z$ and binary outcome $y$, the potential outcomes for unit $i$ are $(y_{i}^{0},y_{i}^{1})$, and these partition the population into four classes or principal strata (Frangakis and Rubin, 2002), which can be modeled as latent outcomes (Page et al., 2015).
@@
Ben-Michael et al. (2024) and Christy and Kowalski (2024) have argued that, once effects are defined (even if only implicitly) at the individual level, it can make sense to consider asymmetric losses using some formalization of the “first do no harm” principle. Analyses using an asymmetric loss function have potential practical implications, because the average utility is no longer the same as the average causal effect.
Consider the case in which $y=1$ is survival and $y=0$ is death. Then the four principal strata correspond to $(y^{0},y^{1})=(1,1)$ (patient survives under either treatment), $(0,0)$ (patient dies under either treatment), $(0,1)$ (treatment saves the patient’s life), and $(1,0)$ (treatment kills the person). In the first two principal strata, the treatment effect is zero. If the negative of the utility in the fourth stratum is taken to be greater than the positive utility of the third stratum, then the resulting decision rule will “prioritize safety over efficacy” (Christy and Kowalski, 2024).
@@
There can be reasonable policy or ethical reasons to prefer low-risk treatments, but we have to be careful, as it is well known that any utility function that depends on unrealized outcomes can lead to incoherent decision recommendations; see, e.g., Bell (1982).
@@
For a very simple example, consider a decision between two options: A, a lottery which yields \$100 with probability 0.6 and \$0 with probability 0.4; and B, which with probability 0.5 yields \$100 and with probability 0.5 returns a lottery which yields \$100 with probability 0.2 and \$0 with probability 0.8. These two lotteries are identical in the distribution of outcomes and thus would receive identical utilities in the classical outcome-based utility framework of von Neumann and Morgenstern (1944). But now consider a utility function in which there is a negative value to uncertainty, which is evaluated at each branch of the tree. For example, suppose the utility of \$100 is set to 1, the utility of \$0 is set to 0, and the existence of uncertainty reduces the utility of any lottery by 10%. Then option A has utility $0.9\,(0.6\cdot 1+0.4\cdot 0)=0.54$, while option B has utility $0.9\,(0.5\cdot 1+0.5\,(0.9\,(0.2\cdot 1+0.8\cdot 0)))=0.9\,(0.5+0.09)=0.531$. Option B here has a lower utility because the utility function pays twice for the uncertainty. This is not to say that such a utility function is “wrong” (or, for that matter, that it is normatively or descriptively right); it just illustrates how it violates the axioms of classical (Neumann-Morgenstern) decision theory.
@@
This example is not just a simple nonlinear utility function for gains and losses (which has its own coherence problems; see section 5 of Gelman, 1998); rather, it violates classical decision theory by depending on unrealized outcomes.
@@
In general, when a decision rule is incoherent, it is possible to construct a scenario in which it yields inappropriate or nonsensical results. Again, that does not mean the rule is never appropriate. Working through simple examples can help us understand its domain of applicability.
@@ 2 Russian roulette
In the case of the “do no harm” principle, we can demonstrate incoherence using a simple example of two independent lotteries:
@@
  • $R_{1/7}$: You live with probability 6/7 and die with probability 1/7,
along with:
  • $R_{0}$: You live.
If you have to choose between $R_{1/6}$ and $R_{1/7}$, the latter should clearly dominate. But, as we shall see, this does not always happen under a regret-based decision rule.
For a decision rule with potential outcomes $(y^{0},y^{1})$, we follow section 3.1 of Christy and Kowalski (2024) and assume the following relative utility function:
$U(z=1\mbox{ relative to }z=0\,|\,y^{0},y^{1})=\left\{\begin{array}[]{rl}0&\mbox{if $y^{1}=y^{0}$}\\ 0.5(y^{1}-y^{0})&\mbox{if $y^{1}>y^{0}$}\\ -(y^{1}-y^{0})&\mbox{if $y^{0}>y^{1}$},\end{array}\right.$ (1)
so that the disutility of a loss is twice as large as the utility of a gain. For binary outcomes $y$, this rule corresponds to:
$U(z=1\mbox{ relative to }z=0\,|\,y^{0},y^{1})=\left\{\begin{array}[]{rl}0&\mbox{if you are sure to live}\\ 0&\mbox{if you are sure to die}\\ +0.5&\mbox{If the treatment would save your life}\\ -1&\mbox{if the treatment were to kill you.}\end{array}\right.$ (2)
First, we separately compare $R_{1/6}$ and $R_{1/7}$ to the status quo of $R_{0}$. Then we consider the scenario in which $R_{1/6}$ is the status quo and we are considering switching to $R_{1/7}$. That is where the counterintuitive result arises.
  • $R_{1/6}$ versus the status quo of $R_{0}$. There is a 5/6 chance that $R_{1/6}$ has no effect and a 1/6 chance that $R_{1/6}$ kills. In the formulation of Christy and Kowalski (2024), the potential outcomes $(y^{0},y^{1})$ are $(1,1)$ for 5/6 of a (large) population and $(1,0)$ for 1/6 of the population. The expected relative utility is $-1/6$.
  • $R_{1/7}$ versus the status quo of $R_{0}$. Similarly, here the expected relative utility is $-1/7$. Unsurprisingly, both $R_{1/6}$ and $R_{1/7}$ have negative utilities compared to $R_{0}$, with $R_{1/6}$ looking worse.
  • $R_{1/7}$ versus the status quo of $R_{1/6}$. We have assumed the outcomes in $R_{1/7}$ and $R_{1/6}$ are independent, so we can easily work out the probabilities of the four potential strata: $(y^{0},y^{1})$ takes on the value $(1,1)$ with probability 30/42, it takes on the value $(0,0)$ with probability 1/42, it is $(1,0)$ with probability 6/42, and it is $(0,1)$ with probability 5/42. Under expression (2), the expected relative utility of $R_{1/7}$ compared to $R_{1/6}$ is thus $0.5(6/42)-5/42=-1/21$.
The result—that, in a head-to-head matchup, we should prefer to stick with $R_{1/6}$—makes no sense. Obviously we would prefer our chance of dying to decrease from 1/6 to 1/7. But, under the asymmetric utility function, you lose more from the 5/42 chance of the new treatment “killing you” than from the 6/42 chance of it “saving your life.”
@@ 3 Moving from deterministic to stochastic potential outcomes
There are two natural ways to resolve this problem and avoid the nonsensical decision recommendation. The first approach would be to move to a classical utility function that depends only on outcomes, in which case one can just evaluate $U(R_{1/6})$, $U(R_{1/7})$, and $U(R_{0})$ directly, and there is no incoherence. This solution is too easy, though, in that it eliminates the “do no harm” principle that motivated the idea in the first place.
A second approach is to preserve the asymmetric utility function but change the nature of the potential outcomes, replacing the deterministic potential outcomes of Neyman (1923) and Rubin (1974) with stochastic potential outcomes as in Greenland (1987) and Vanderweele and Robins (2012). In the Russian roulette example, this entails replacing the binary outcome $y^{z}$ with a binary random variable whose expectation is the probability of living under treatment $z$. We then generalize the utility function in Equation 1 to allow stochastic potential outcomes:
$U(z=1\mbox{ relative to }z=0\,|\,y^{0},y^{1})=\left\{\begin{array}[]{rl}0&\mbox{if $U(y^{1})=U(y^{0})$}\\ 0.5(U(y^{1})-U(y^{0}))&\mbox{if $U(y^{1})>U(y^{0})$}\\ -(U(y^{1})-U(y^{0}))&\mbox{if $U(y^{0})>U(y^{1})$.}\end{array}\right.$ (3)
In the problem we are considering, the $y^{z}$’s are binary random variables, and we set $U(y^{z})=E(y^{z})$. What is relevant in (3) is that the utilities of $y^{0}$ and $y^{1}$ are evaluated individually, and then the asymmetric utility for the decision is applied to these separate utilities. This follows the same principles as (1), with the only change being that the potential outcomes $y^{0}$ and $y^{1}$ are random variables.
Here are the three decision problems under probabilistic potential outcomes:
  • $R_{1/6}$ versus the status quo of $R_{0}$. Here, $U(y^{1})=5/6$ and $U(y^{0})=1$ for everyone in the population, and so the decision will certainly reduce your probability of survival. The expected relative utility is $-1/6$.
  • $R_{1/7}$ versus the status quo of $R_{0}$. Similarly, here the expected relative utility is $-1/7$.
  • $R_{1/7}$ versus the status quo of $R_{1/6}$. Here, $U(y^{1})=6/7$ and $U(y^{0})=1/6$ for everyone in the population, so the decision will certainly increase your probability of survival. The expected relative utility is $0.5(6/7-5/6)=1/84$, which is positive, as it should be.
Paradoxically, when we use stochastic potential outcomes in this example, the utility functions become deterministic. In general, this will not necessarily be the case. For example, consider a hypothetical medical condition where the survival rate is evenly distributed between 20% and 50% (depending current health status), and there is a treatment that only works on 20% of the population, but when it does work, it halves the death rate. In that case, one can set up a mixture model for potential outcomes in the population, where $y^{0}$ is uniformly distributed in the range $(0.2,0.5)$ and $y^{1}=y^{0}$ for 80% of the population and $y^{1}=1-(1-y^{0})/2$ for the remaining 20%. Such a model could be elaborated by allowing the mixture probability to depend on $y^{0}$ and by letting the entire distribution depend on patient-level variables $x$.
@@ 4 Stochastic potential outcomes and the stable unit treatment value assumption
Stochastic models for potential outcomes require a generalization of SUTVA (the stable unit treatment value assumption; Rubin, 1980). The issue here arises not with the first assumption of SUTVA, the familiar rule of no interference within units, but rather with the lesser-known second assumption, which Rubin (2005) expresses as “there are no hidden versions of treatments; no matter how unit $i$ received treatment 1, the outcome that would be observed would be $Y_{i}(1)$ and similarly for treatment 0.” In the Russian roulette example there are no “hidden versions” of treatments in the usual sense of the phrase—you spin the cylinder and fire the shot; it’s the same treatment every time—but the potential outcomes are not deterministic.
@@
That is fine. SUTVA is not a requirement for causal inference; it is just a particular set of assumptions, appropriate in some settings but not for others. Just as we lift the no-interference-within-units assumption when modeling spillover effects, we can lift the second assumption of SUTVA when considering treatments that yield stochastic potential outcomes. There are many problems in biomedical and social research where outcomes are inherently noisy, so that such stochastic causal models make sense, even for treatments that are applied the same way to each unit.
@@
Stochastic potential outcomes is not a new idea—Neyman et al. (1935) refer to them as “technical errors.” In classical decision theory, they can be thought of as unobserved latent variables, so that any stochastic potential outcome model can be rewritten as a distribution over deterministic values, in the same way that logistic regression can be viewed as an integration over logistically-distributed latent variables. Vanderweele and Robins (2012) explain how a stochastic potential outcomes model can facilitate a convenient framework for modeling multiple causes.
@@
This modeling choice is reminiscent of the distinction in classical statistical theory between parameters, which are conditioned on when evaluating inferential methods, and latent data or predictions, which are averaged over. For example, classical statistics distinguishes between unbiased estimation of parameters $\theta$ (the condition $\mbox{E}(\hat{\theta}(y)|\theta)=\theta$) and unbiased prediction of latent data $u$ (the condition $\mbox{E}(\hat{u}(y)|\theta)=\mbox{E}(u|\theta)$). In Bayesian inference, there is no such distinction—all parameters are considered to be a form of latent data and are given probability distributions. This is consistent with the property of the Neumann-Morgenstern framework (which, confusingly, we call classical decision theory although it is fully Bayesian) that there is no fundamental distinction between deterministic and stochastic potential outcomes.
@@
Even setting aside the problem of utilities that depend on unrealized potential outcomes, we have found stochastic potential outcomes to be useful as a way of partitioning into unit-specific and unexplained variation (Gelman, 2021). More formally, one can imagine hypothetical multiple applications of the treatment to the same unit: in the Russian roulette example, different trials could yield different results on the same person, whereas if the potential outcome depends entirely on person-level characteristics, it would be the same every time. As discussed in the next section, the partition of variation depends on the extent to which the potential outcomes depend on immutable (if unobserved) unit-level characteristics, which in turn depends on the application.

### v2

@@ Abstract
It has been proposed in medical decision analysis to express the “first do no harm” principle as an asymmetric utility function in which the loss from killing a patient would count more than the gain from saving a life. Such a utility depends on unrealized potential outcomes, and we show how this yields a paradoxical decision recommendation in a simple hypothetical example involving games of Russian roulette. The problem is resolved if we abandon the stable unit treatment value assumption (SUTVA) and allow the potential outcomes to be random variables. This leads us to conclude that, if you are interested in this sort of asymmetric utility function, you need to move to the stochastic potential outcome framework. We discuss the implications of the choice of parameterization in this setting.
1 Introduction
1.1 Utility functions for potential outcomes in causal inference
@@
Causal inference has traditionally focused on identification of average effects. In recent decades, there has been increasing interest in considering different forms of the average causal effect [14], flexible models for individual effects [13, 24], and understanding how effects vary [5].
@@
Individual causal effects can never be precisely estimated (without multiple measurements on the same units and strong assumptions about crossover effects), but they can be defined theoretically as latent variables using the potential-outcomes framework [16, 19]. In the simplest case of binary treatment $z$ and binary outcome $y$, the potential outcomes for unit $i$ are $(y_{i}^{0},y_{i}^{1})$, and these partition the population into four classes or principal strata [6], which can be modeled as latent outcomes [18].
@@
[2] and [3] have argued that, once effects are defined (even if only implicitly) at the individual level, it can make sense to consider asymmetric losses using some formalization of the “first do no harm” principle. Analyses using an asymmetric loss function have potential practical implications, because the average utility is no longer the same as the average causal effect. Utility functions over potential outcomes have also been considered by [4] and [12].
1.2 Relation to classical utility theory
@@
There can be reasonable policy or ethical reasons to prefer low-risk treatments, but we have to be careful, as it is well known that any utility function that depends on unrealized outcomes can lead to incoherent decision recommendations; see, e.g., [1].
@@
For a simple example of an incoherent decision rule when moving beyond the classical outcome-based utility framework of [15], consider a decision between two options: A, a lottery which yields \$100 with probability 0.6 and \$0 with probability 0.4; and B, which with probability 0.5 yields \$100 and with probability 0.5 returns a lottery which yields \$100 with probability 0.2 and \$0 with probability 0.8. These two lotteries are identical in the distribution of outcomes and thus would receive identical utilities in the classical Neumann-Morgenstern framework. But now consider a utility function in which there is a negative value to uncertainty, which is evaluated at each branch of the tree. For example, suppose the utility of \$100 is set to 1, the utility of \$0 is set to 0, and the existence of uncertainty reduces the utility of any lottery by 10%. Then option A has utility $0.9\,(0.6\cdot 1+0.4\cdot 0)=0.54$, while option B has utility $0.9\,(0.5\cdot 1+0.5\,(0.9\,(0.2\cdot 1+0.8\cdot 0)))=0.9\,(0.5+0.09)=0.531$. Option B here has a lower utility because the utility function pays twice for the uncertainty. This is not to say that such a utility function is “wrong” (or, for that matter, that it is normatively or descriptively right); it just illustrates how it violates the axioms of classical (Neumann-Morgenstern) decision theory.
@@
This example is not just a simple nonlinear utility function for gains and losses (which has its own coherence problems; see section 5 of [7]); rather, it violates classical decision theory by depending on unrealized outcomes.
@@
When a decision rule is incoherent, it should be possible to construct a “money pump” or some other scenario that yields inappropriate or nonsensical decision recommendations. Again, that does not mean the rule is never appropriate. Working through simple examples can help us understand its domain of applicability.
1.3 Expressing the “first do no harm” principle as an asymmetric utility function
Suppose we want to decide between two possible treatments, where $z=0$ is the “control” or status quo and $z=1$ is a proposed new treatment. We follow [3, Section 3.1] in using a decision rule based on potential outcomes $(y^{0},y^{1})$, where $y^{0}$ is the outcome under $z=0$ and $y^{1}$ under $z=1$. Specifically, we assume the following relative utility function:
$U(z=1\mbox{ relative to }z=0\,|\,y^{0},y^{1})=\left\{\begin{array}[]{rl}0&\mbox{if $U(y^{1})=U(y^{0})$}\\ 0.5(U(y^{1})-U(y^{0}))&\mbox{if $U(y^{1})>U(y^{0})$}\\ -(U(y^{1})-U(y^{0}))&\mbox{if $U(y^{0})>U(y^{1})$,}\end{array}\right$. (1)
so that the disutility of a loss is twice as large as the utility of a gain. We introduce the notation “relative to” to highlight the asymmetry of this utility function designed to favor the status quo.
Consider the case in which $y=1$ is survival and $y=0$ is death. Then the four principal strata correspond to $(y^{0},y^{1})=(1,1)$ (patient survives under either treatment), $(0,0)$ (patient dies under either treatment), $(0,1)$ (treatment saves the patient’s life), and $(1,0)$ (treatment kills the person). In the first two principal strata, the treatment effect is zero. If the negative of the utility in the fourth stratum is taken to be greater than the positive utility of the third stratum, then the resulting decision rule will “prioritize safety over efficacy” [3]. Without loss of generality, we can write $U(0)=0$ and $U(1)=1$, and equation (1) becomes,
[…]
@@ 2 Russian roulette
In the case of the “do no harm” principle, we can demonstrate the incoherence of the above asymmetric utility function using a simple example of two independent lotteries:
@@
  • $R_{1/7}$: You live with probability 6/7 and die with probability 1/7.
It is a clear improvement to switch from the status quo “control” of $R_{1/6}$ to the “treatment,” $R_{1/7}$. But, as we shall see, this is not necessarily recommended under a regret-based decision rule.
We assume the outcomes in $R_{1/6}$ and $R_{1/7}$ are independent, so we can easily work out the probabilities of the four potential strata: $(y^{0},y^{1})$ takes on the value $(1,1)$ with probability 30/42, it takes on the value $(0,0)$ with probability 1/42, it is $(1,0)$ with probability 5/42, and it is $(0,1)$ with probability 6/42. Under expression (2), the expected relative utility of $R_{1/7}$ compared to $R_{1/6}$ is thus 0.5(6/42)5/42=1/210.5(6/42)-5/42=-1/21. Here the potential outcomes are fixed but unknown quantities, and the probabilities correspond to uncertainty about the states of the guns or, equivalently, as averages over a hypothetical superpopulation of plays.
This result—that we should stick with $R_{1/6}$—makes no sense. Obviously we would prefer our chance of dying to decrease from 1/6 to 1/7. But, under the asymmetric utility function, you lose more from the 5/42 chance of the new treatment “killing you” than you gain from the 6/42 chance of it “saving your life.”
We can also work out the above algebra under an alternative formulation. Let $\phi$ be a random variable that equals 1 if the chamber is loaded with a bullet when the player draws the trigger and 0 otherwise. We then have the deterministic potential outcomes $y(\phi=0)=1$ (alive) and $y(\phi=1)=0$ (dead). In this parameterization, all randomness is induced by the distribution over $\phi$ induced by the treatment $z$. Different lotteries $R_{p}$ correspond to different distributions of the intermediate treatment, $\phi$. In this situation, $(y^{0},y^{1})=(y(\phi),y(\phi^{\prime}))$ with independent random variables $\phi\sim\mbox{Bernoulli}(1/6)$ and $\phi^{\prime}\sim\mbox{Bernoulli}(1/7)$. Expression (1) then becomes,
$\mbox{E}_{\phi,\phi^{\prime}}\left(U(R_{1/7}\mbox{ relative to }R_{1/6}\,|\,y^{0},y^{1})\right)\,$ $=\,\mbox{E}_{\phi,\phi^{\prime}}\left(0.5\cdot 1_{y^{1}>y^{0}}-1_{y^{0}>y^{1}}\right)$
= $=\,0.5\cdot\mbox{Pr}(\phi>\phi^{\prime})-\mbox{Pr}(\phi^{\prime}>\phi)$
$=0.5(1/7)-5/42=-1/21,$
and again the asymmetric decision rule yields an unacceptable conclusion.
@@ 3 Moving from deterministic to stochastic potential outcomes
There are two natural ways to resolve this problem and avoid the nonsensical decision recommendation. The first approach would be to move to a classical utility function that depends only on outcomes, in which case one can just evaluate $U(R_{1/6})$ and $U(R_{1/7})$, and there is no incoherence. This solution is too easy, though, in that it eliminates the “do no harm” principle that motivated the idea in the first place.
A second approach preserves the asymmetric utility but changes the nature of the potential outcomes, replacing the deterministic potential outcomes of [16] and [19] with stochastic potential outcomes as in [11] and [23].
To adapt (1) to allow stochastic potential outcomes $y^{z}$ (where $z=0$ or 1, corresponding to the $R_{1/6}$ or $R_{1/7}$, respectively), the utility of a potential outcome now needs to take its variability into account. In the Russian roulette example, we replace each binary outcome $y^{z}$ with a binary random variable whose expectation is the probability of survival under treatment $z$. The Russian roulette treatment thus includes the spinning of the cylinder as well as the pulling of the trigger, and the assumption is that the outcome of the spin is random and is not predictable by any characteristics of the individual. It’s natural to average over this level of variability when evaluating the utilities $U(y^{z})=E(y^{z})$, that is, to define the utility as the probability of survival. What is relevant here is that the utilities of $y^{0}$ and $y^{1}$ are evaluated individually, and then the asymmetric utility for the decision is applied to these separate utilities. This follows the same principles as in the previous section, with the only change being that the potential outcomes $y^{0}$ and $y^{1}$ are random variables.
When evaluating $R_{1/7}$ versus the status quo of $R_{1/6}$, the utilities are $U(y^{1})=6/7$ and $U(y^{0})=1/6$ for everyone in the population, so the decision will certainly increase survival probability. The expected relative utility is 0.5(6/75/6)=1/840.5\,(6/7-5/6)=1/84, which is positive, as it should be.
When switching from deterministic to stochastic potential outcomes the order between evaluating utilities and integration over uncertainty changes. In the deterministic case, we first evaluate the utilities and then integrate over the possible treatment assignments under the policies $R_{p}$. When using stochastic potential outcomes, the utilities are evaluated on random variables, collapsing the uncertainty in a unit’s outcome under a specific treatment.
@@ 4 Stochastic potential outcomes and the stable unit treatment value assumption
Stochastic models for potential outcomes require a generalization of SUTVA (the stable unit treatment value assumption; [20]). The issue here arises not with the first assumption of SUTVA, the familiar rule of no interference within units, but rather with the lesser-known second assumption, which [21] expresses as “there are no hidden versions of treatments; no matter how unit $i$ received treatment 1, the outcome that would be observed would be $Y_{i}(1)$ and similarly for treatment 0.” In the Russian roulette example, there are no “hidden versions” of treatments in the usual sense of the phrase—you spin the cylinder and fire the shot; it’s the same treatment every time—but the potential outcomes are not deterministic.
@@
That is fine. SUTVA is not a requirement for causal inference; it is just a set of assumptions that are appropriate in some settings but not for others. Just as we lift the no-interference-within-units assumption when modeling spillover effects, we can lift the second assumption of SUTVA when considering treatments that yield stochastic potential outcomes. There are many problems in biomedical and social research where outcomes are inherently noisy, so that such stochastic causal models make sense, even for treatments that are applied the same way to each unit.
@@
Stochastic potential outcomes is not a new idea—[17] refers to them as “technical errors.” In classical decision theory, they can be thought of as unobserved latent variables, so that any stochastic potential outcome model can be rewritten as a distribution over deterministic values, in the same way that logistic regression can be viewed as an integration over logistically-distributed latent variables. [23] explain how a stochastic potential outcomes model can facilitate a convenient framework for modeling multiple causes.
@@
This modeling choice is reminiscent of the distinction in classical statistical theory between parameters, which are conditioned on when evaluating inferential methods, and latent data or predictions, which are averaged over. For example, classical statistics distinguishes between unbiased estimation of parameters $\theta$ (the condition $\mbox{E}(\hat{\theta}(y)|\theta)=\theta$) and unbiased prediction of latent data $u$ (the condition $\mbox{E}(\hat{u}(y)|\theta)=\mbox{E}(u|\theta)$). In Bayesian inference, there is no such distinction—all parameters are considered to be a form of latent data and are given probability distributions. This is consistent with the property of the Neumann-Morgenstern framework (which, confusingly, is often called classical decision theory although it is fully Bayesian) that there is no fundamental distinction between deterministic and stochastic potential outcomes.
@@
Even setting aside the problem of utilities that depend on unrealized potential outcomes, we have found stochastic potential outcomes to be useful as a way of partitioning into unit-specific and unexplained variation [8]. More formally, one can imagine hypothetical multiple applications of the treatment to the same unit: in the Russian roulette example, different trials could yield different results on the same person, whereas if the potential outcome depends entirely on person-level characteristics, it would be the same every time. As discussed in the next section, the partition of variation depends on the extent to which the potential outcomes depend on immutable (if unobserved) unit-level characteristics, which in turn depend on the application.